Half a pound of gas undergoes a reversible non-flow process with the pressure remaining constant at 50 psia while the volume increases from 4 ft' to 8ft. Find the work done, Btu/lb


1
Expert's answer
2021-09-14T17:59:02-0400

We only have to substitute the solution for the integral of the work when the pressure remains constant:


"W=\int_{V_1}^{V_2}pdV =p{\Delta V}=p(V_2-V_1)
\\ \text{ }
\\ \text{then, we substitute and find the value for W:}
\\W=\dfrac{(50\,psia)(8-4)\,ft^3}{0.5 \,lb}
\\ \text{ }
\\ \text{finally, we use the relation between the variables to find the value for W in Btu/lb:}
\\W=400\,\dfrac{psia\cdot ft^3}{lb} \times\cfrac{1\,Btu}{5.40362\,psia\cdot ft^3}
\\W=74.0244\, \cfrac{Btu}{lb}"


In conclusion, we find that the work done is equal to W=74.0244 But/lb.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.